Circular No. 7483

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000di IN IC 1637 S. Nakano, Sumoto, Japan, reports the discovery by M. Aoki, Tsukioka,

Toyama, of an apparent supernova (mag 17.3) on a CCD frame (limiting mag 18.5) taken on Aug. 23.708 UT with his automated supernova survey system employing a 0.43-m f/6.0 reflector. SN 2000di (still at mag 17.3) was confirmed on seven frames taken on Aug. 24.71 (limiting mag 17.5). The new object is quite close to an H II region; the supernova is located at α = 1^{h}11^{m}02^{s}.01, δ = −30^{o}26′31′′.2 (equinox 2000.0), which is 12′′.3 east and 11′′.5 south of the nucleus of IC 1637.

A. V. Filippenko and W. D. Li, University of California, Berkeley (UCB), report that inspection of a CCD spectrum (range 330–1000 nm), obtained on Aug. 27 UT with the Shane 3-m reflector at Lick Observatory, reveals SN 2000di to be a type-II supernova; the spectrum resembles that of SN 2000dc (see below).

SUPERNOVAE 2000dc, 2000dd, AND 2000df D. C. Leonard, W. D. Li, and A. V. Filippenko, UCB, report that in-

spection of a CCD spectrum (range 320–1000 nm) of SN 2000df (cf. IAUC 7478), obtained as above on Aug. 26 UT, reveals that it is a type-Ia supernova, ∼ 2 to 3 weeks past maximum. Similarly, a spectrum of SN 2000dc (IAUC 7476) shows it to be of type II, with well-developed Hβ and Fe II (501.8, 516.9 nm) P-Cyg absorption. Prominent, broad Hα emission is present, but the absorption profile is peculiar, showing two relatively weak absorption minima at 16 200 and 8500 km/s, similar to features seen in SN 1999gk (IAUC 7332) and SN 1998S (Leonard et al. 2000, Ap.J. 536,

239) early in their development. It is possible that the higher-velocity component is actually due to Si II 635.5-nm, as this identification brings the feature’s expansion velocity (6500 km/s) into agreement with that derived from the other weak metal lines. Filippenko and Li add that inspection of a CCD spectrogram (obtained as above) shows that SN 2000dd (IAUC 7477) is of type Ia, roughly 2 weeks past maximum brightness.

NO SUPERNOVA 2000dh Filippenko and Li report that a CCD spectrum (obtained at Lick as

noted above) indicates that SN 2000dh (IAUC 7482) is actually an M-type star in our Galaxy: strong TiO absorption features are visible in the nearinfrared. Some such stars are known to exhibit considerable variability.

2000 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green